OpenStudy (anonymous):

how do I complete the square? y= x^2 - 6x + 7

OpenStudy (anonymous):

$$(x-3)^2-2=x^2-6x+7$$

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

$$(x-a)^2=x^2-2ax+a^2$$

OpenStudy (anonymous):

Subtract/add to the left hand side until the constant on the right is what you want

OpenStudy (anonymous):

\[x^2 - 6x + 7 =x^2 - 6x +9-9 +7=x^2 - 6x +3^2-2=(x-3)^2-(\sqrt{2})^2\]

OpenStudy (anonymous):

hmm

OpenStudy (blurbendy):

do you understand?

OpenStudy (anonymous):

not really D:

OpenStudy (blurbendy):

http://www.purplemath.com/modules/sqrquad.htm look at the example in the box

OpenStudy (anonymous):

Dear Laurenn_215, in order to complete the square of the given function check the coeficient of x i.e.(- 6x) it is -6 next find (-6)/2=-3 find the square of (-3) i.e. (-3)^2=9 Now add and subtract 9 in the given function so it will be \[x^2 - 6x + 7 = x^2 - 6x + 9-9+7 = x^2 - 2\times x \times 3 + 9-2\] \[= x^2 - 2\times x \times 3 + 3^2-2\] Now this part i.e. \[x^2 - 2\times x \times 3 + 3^2\] is written in the form of \[a^2 - 2\times a \times b + b^2 = (a-b)^2\] so x^2 - 2times x times 3 + 3^2-2 can be expressed as \[= (x-3)^2-2\]

OpenStudy (anonymous):

thanks everyone